OpenStudy (anonymous):

Given: X/5 + 9 =11 Prove: x=10 x/5 + 9 = 11 a.___________ x/5 = 2 b.____________ x=10 c.____________

OpenStudy (anonymous):

A. a. Given b. Subtraction Property of Equality c. Division Property of Equality B. a. Given b. Subtraction Property of Equality c. Multiplication Property of Equality C. a. Given b. Addition Property of Equality c. Multiplication Property of Equality D. a. Given b. Addition Property of Equality c. Division Property of Equality

OpenStudy (texaschic101):

1. given 2. subtraction property -- what you subtract on one side, you subtract on the other side x/5 + 9 = 11 x/5 - 9 = 11 - 9 combine like terms x/5 = 2 3. multiplication property of equality -- what you multiply on one side, you multiply on the other x/5 (5) = 2(5) --- the 5's on the left cancel out x = 10

OpenStudy (texaschic101):

oops...step 2 x/5 + 9 - 9 = 11 - 9

OpenStudy (anonymous):

@texaschic101 muchas gracias

OpenStudy (texaschic101):

denada :)